Suppose you borrow ​$950 for a term of three years at simple interest and 5.65​% APR. Determine the total​ (principal plus​ interest) you must pay back on the loan.

The total​ (principal plus​ interest) you must pay back on the loan is

​$_.
Answer #1

Simple interest=Principal*Interest rate*Time period

which is equal to

=(950 * 3 * 5.65%) = $ 161.025

Hence total to be paid=Simple interest+Principal

=( 161.025 + 950 )

which is equal to

=$1111.02 (Approx).

.